OpenStudy (anonymous):

I need help understanding integration by substitution: 63/(9x+2)^8

OpenStudy (phi):

if you let u = 9x+2 what is du ?

OpenStudy (xapproachesinfinity):

This is the as saying f(x) =9x+2 What is f'(x) Just a difference of differentials but you don't need to worry about thst

OpenStudy (anonymous):

du=9 right?

OpenStudy (xapproachesinfinity):

Yes just one thing you missed Du=9dx

OpenStudy (phi):

no du/dx =9 but if you "take the derivative" u= 9x you differentiate the variables, in this case the u and the x like this du = 9 dx

OpenStudy (phi):

and once you have du = 9 dx then you also have \[ dx = \frac{1}{9} du \]

OpenStudy (phi):

now do your variable substitution replace 9x+2 with u replace dx with 1/9 du

OpenStudy (phi):

and if we have limits, we would also change the limits.

OpenStudy (phi):

and because it is easy to integrate a power, I would use the -8 exponent and get rid of the fraction

OpenStudy (anonymous):

So would it be: \[63u ^{-8}*1/9 du\] ??

OpenStudy (phi):

yes and the problem is \[ 7 \int u^{-8} du \] which you can do , right?

OpenStudy (xapproachesinfinity):

Yes

OpenStudy (phi):

once you integrate, replace u with 9x+2 to put the answer in terms of x

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{63}{(9x+2)^8}~dx}\) \(\large\color{slate}{\displaystyle u=9x+2}\) \(\large\color{slate}{\displaystyle du=(9x+2)'~\cdot dx~~~\rightarrow~~~du=9~dx}\) (you already have a 9 and dx to replace that by du, just need to re-write it. \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{7\cdot 9}{(9x+2)^8}~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{7}{(9x+2)^8}~(9\cdot dx)}\) substitution:: \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{7}{(u)^8}~(du)}\) and then all you have left:: \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}7~u^{-8}~du}\)

OpenStudy (solomonzelman):

(APPLY THE POWER RULE, AND don't forget to substitute the x back for u.)

OpenStudy (solomonzelman):

igtg....

OpenStudy (anonymous):

Ok, I got it! Thank you all!